OpenStudy (starlover4556):

Help Please !!!!! OK so I just graphed using slope intercept form and on the next question it says why don't the lines start at the origin. Now the lines I have graphed don't start at the origin but the problem is I no clue why they don't. Is there a certain reason why they might not Please help.

OpenStudy (g3n3sis08):

They are asking why it doesn't.. what's the equation?

OpenStudy (starlover4556):

it is y=25+10x and y=18+15x

OpenStudy (g3n3sis08):

Whats the first question please?

OpenStudy (starlover4556):

Two cell phone companies have different rate plans. Runfast has monthly charges $25 plus $10 per gig of data. B A &D’s monthly charge is $18 plus 15 per gig of data. Your task is to determine under what circumstances each company has the better pricing. You will derive and solve a set of equations, graph those equations, and evaluate the meanings for these events. Don’t worry; I have a series of questions to guide you through the process. Good luck! 1. Determine the equation for the monthly charges for Runfast and B A&D.

OpenStudy (starlover4556):

Number 2 said solve this system of equations using the substitution method i did that and i got x=7/5 and y=39

OpenStudy (welshfella):

they dont start at the origin becuase of the montgly standing charges. You pay this whether you use the phone or not.

OpenStudy (welshfella):

your solutions are correct

OpenStudy (welshfella):

those numbers correspond to where the graphs of the equations intersect.

OpenStudy (starlover4556):

ok good i thought I might have done that wrong

OpenStudy (starlover4556):

this is what my teacher wanted us to use to graph using slope intercept form

OpenStudy (welshfella):